OpenStudy (anonymous):

Need help please! Brennan has been playing a game where he can create towns and help his empire expand. Each town he has allows him to create 1.15 times as many villagers. The game gave Brennan 5 villagers to start with. Help Brennan expand his empire by solving for how many villagers he can create with 15 towns. Then explain to Brennan how to create an equation to predict the number of villagers for any number of towns. Show your work and use complete sentences.

OpenStudy (anonymous):

Just need someone to walk me through what I should do.

OpenStudy (alivejeremy):

still need help

OpenStudy (anonymous):

Yeah

OpenStudy (alivejeremy):

ok i got u

OpenStudy (alivejeremy):

so are u trying to work this out

OpenStudy (sparklyme):

y is how many villagers he will get with each town. x is the number of the town example: town # 7 y=1.15x7+5 this would mean that y=8.05

OpenStudy (alivejeremy):

nice job

OpenStudy (alivejeremy):

but i could have done tht

OpenStudy (sparklyme):

@alivejeremy Me?

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

yea u

OpenStudy (alivejeremy):

@sparklyme

OpenStudy (anonymous):

@sparklyme You just got that off of Yahoo Answers, and anyways I need to know how to do this -- not the answer.

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

Yahoo

OpenStudy (alivejeremy):

lol

OpenStudy (anonymous):

@Michele_Laino Can you help me out?

OpenStudy (alivejeremy):

sure

OpenStudy (alivejeremy):

You have a geometric progression with $a = 5$ and $r = 1.15$. The sum of the first $n$ terms of this progression is $ s_n = 5\left(\frac{1-(1.15)^n}{1-1.15}\right)$. The answer to the first part of the question is $ floor(s_15 = 5\left( \frac{1-(1.15)^{15}}{1-1.15}\right)) = floor(47.58) = 47$ villagers. To find the answer to the second part, you know that there are $s_n$ villagers. To find the number of towns you have to assume hat the maximum number of villagers has been created and rearrange $ s_n = 5\left( \frac{1-(1.15)^n}{1-1.15}\right)$ to make $n$ the subject, a follows: 5(1−(1.15)n)=(1−1.15)Sn=−0.15Sn5−5(1−(1.15)n)=−0.15Sn5(1.15)n=5+0.15Sn(1.15)n=5+0.15Sn5Sn=log1.155+0.15x5

OpenStudy (alivejeremy):

u get it

OpenStudy (alivejeremy):

i tryed to break it down

OpenStudy (anonymous):

That's not helping, don't copy and paste answers from other people's responses.

OpenStudy (alivejeremy):

no i didn't someone help me wit tht question before

OpenStudy (alivejeremy):

K12

OpenStudy (anonymous):

It's obviously someone else's answer, I just looked it up.

OpenStudy (alivejeremy):

ik u did

OpenStudy (alivejeremy):

but tht person help me to

OpenStudy (alivejeremy):

As it compounds, we can clearly see that every time he creates a town, the number of people increase by 1.15 times. so he creates a total of 15 towns, and he has 5 villagers to start with. So you would have to multiply it like ((5*1.15)) for the first village, and hten (5*1.15)*1.15 for the second village and so on. Therefore for 15 villages, it has to be (5*(1.15^15)), which if you have a calculator, is 40.6853081458. :P Therefore its quite obvious that to create more villages, the equation would have to be (5*(1.15^n)), where n is the number of villages being made one after the other. You have a geometric progression with $a = 5$ and $r = 1.15$. The sum of the first $n$ terms of this progression is $ s_n = 5\left(\frac{1-(1.15)^n}{1-1.15}\right)$. The answer to the first part of the question is $ floor(s_15 = 5\left( \frac{1-(1.15)^{15}}{1-1.15}\right)) = floor(47.58) = 47$ villagers. To find the answer to the second part, you know that there are $s_n$ villagers. To find the number of towns you have to assume hat the maximum number of villagers has been created and rearrange $ s_n = 5\left( \frac{1-(1.15)^n}{1-1.15}\right)$ to make $n$ the subject. basically the equation is a prediction of how many villagers will be in each town. Hope this helps, have a nice day. Tag me on anymore you need help with.

OpenStudy (michele_laino):

hint: if we start wit \(a_1=5\) villagers, then after the first town we have \(5 \cdot 1.15=5.75\) villagers, namely \(a_2=1.15a_1\) villagers

OpenStudy (alivejeremy):

i copy and paste tht one tho

OpenStudy (michele_laino):

at the next step, we have \(a_3=1.15 a_2\) villagers

OpenStudy (alivejeremy):

@Michele_Laino dude u always give hints

OpenStudy (alivejeremy):

he want u to break it down

OpenStudy (anonymous):

I'm still not really understanding, am I supposed to plug in the first term and ratio of 1.15 into a geometric formula?

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

no i didn't someone help me wit tht question before

OpenStudy (michele_laino):

hint: in other words, we get the subsequent sequence: \(a_1,\;a_2=1.15a_1,\;a_3=1.15a_2,\;...\) and so on...

OpenStudy (michele_laino):

of course \(a_1=5\)

OpenStudy (anonymous):

So 15 towns is a_1 [or 5] multiplied by 1.15 15 times?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (michele_laino):

it is a geometric sequence

OpenStudy (anonymous):

Which would be 86.25? So 86.25 villagers can be created with 15 towns

OpenStudy (michele_laino):

the first term is \(a_1\), whereas the last term, after 15 towns, is \(a_{16}\), so we can write this: \[\Large {a_{16}} = {a_1} \cdot {1.15^{15}}\]

OpenStudy (michele_laino):

since I have applied the general formula: \[\Large {a_n} = {a_1}{q^{n - 1}}\] and \(q=1.15\)

OpenStudy (anonymous):

So an equation to find the nth amount of villagers for any amount of towns would be a(n) = a*1.15^n-1?

OpenStudy (michele_laino):

yes! with \(n=16\)

OpenStudy (anonymous):

Ohh okay. So the first part of the question can be answered: 86.25 villagers can be made from 15 towns because a16=a1*1.15^15

OpenStudy (anonymous):

and the second part can be written as: a(n)=a*1.15^(16-1) is the expression that can be used to find the nth term in this geometric sequence... or something like that?

OpenStudy (michele_laino):

we can write this: \[\huge {a_n} = {a_1}\cdot {1.15 ^{(n - 1)}}\]

OpenStudy (michele_laino):

with \(a_1=5\)

OpenStudy (anonymous):

Okay, I think I understand it now, thanks

OpenStudy (michele_laino):

:)